Let f:R2→R2 be defined by f(x,y)=(-y,-x)

i) show that f is linear

ii)Determine a basis for the kernel of f and the nullity of f

iii) Determine the basis for the range of f and the rank of f

iv) Determine whether f is invertible or not



1
Expert's answer
2021-09-09T16:32:01-0400

i) Let f:R2→R2 be defined by f(x,y)=(-y,-x)

Let "\beta =\{ (1,0)= e_1, (0,1)= e_2 \}" be a standard ordered pair of R2

Then f(0,1) = (1,0)= 1(1,0)+0(0,1)

"[F]_{\beta}=\begin{bmatrix}
 0 & 1\\
 1 & 0
\end{bmatrix}= A(ay)"

"A(x)= |A-xI| = \begin{vmatrix}
 -x & 1 \\
 1 & -x
\end{vmatrix}= x^2-1" Hence linear

ii)

"F(x)= Ax\\
A= \begin{pmatrix}
 1 & 1 \\
 1 & -1
\end{pmatrix}"

"dom = R^2, co-dom = R^2\\
dim \space dom = 2, dim \space co-dom = 2\\
Ax= 0 \implies \begin{pmatrix}
 1 & 1 \\
 1 & -1
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix} = \begin{pmatrix}
 0 \\
 0
\end{pmatrix}\\
\implies \begin{pmatrix}
 x+y\\
 x-y
\end{pmatrix}= \begin{pmatrix}
 0 \\
 0
\end{pmatrix}\\
x+y=0\\
x-y=0\\
\implies x=0; y=0\\
\therefore Kernel \space F = \{(0,0)\} \\
Nullity =0\\"


iii)

"dom = R^2, co-dom = R^2\\
dim \space dom = 2, dim \space co-dom = 2\\
Ax= 0 \implies \begin{pmatrix}
 1 & 1 \\
 1 & -1
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix} = \begin{pmatrix}
 0 \\
 0
\end{pmatrix}\\
\implies \begin{pmatrix}
 x+y\\
 x-y
\end{pmatrix}= \begin{pmatrix}
 0 \\
 0
\end{pmatrix}\\
x+y=0\\
x-y=0\\
\implies x=0; y=0\\
\therefore Range \space F = R^2\\
\implies Range = \{(1,0), (0,1)\}"


iv)

"For \space y \epsilon R" to find "x \epsilon R" such that

"f(x)=y\\
f(x,y)=(-y,-x)"

This shows that f is invertible